# Under-coverage in high-statistics counting experiments with finite MC samples

### v2

@@ 3.2 Impact of MC fluctuations on Hessian uncertainties
The results presented in section 3.1.1 assume that the model function ${\bf f}$ is known exactly. Suppose instead that ${\bf f}$ is determined from MC samples consisting of finite numbers of events, thus resulting in perturbed estimators $\hat{\bf b}={\bf b}+\boldsymbol{\beta}$ and $\hat{\bf A}$, with the latter giving rise to a perturbed matrix $\hat{\bf U}$, defined as in eq. (30) with the replacement ${\bf A}\to\hat{\bf A}$. Since $\hat{\bf U}$ must also be idempotent, it admits an expansion in terms of a new basis of eigenvectors $\hat{\bf u}_{j}={\bf u}_{j}+\boldsymbol{\nu}_{j}$. By expanding $S$ around its true value, we thus obtain an estimator $\hat{S}$ given by
where the dots stand for terms that are of higher-order in the perturbations. We now take the expectation value of the right-hand side of eq. (3.2) in the space of MC predictions, that is the mean value of $\hat{S}$ if infinitely many identical replicas of the MC templates could be tested, obtaining:
@@
$-2\sum_{j=1}^{p}\left({\rm Tr}\left[\langle\boldsymbol{\nu}_{j}\boldsymbol{\beta}^{T}\rangle\,{\bf u}_{j}{\bf b}^{T}\right]+{\rm Tr}\left[\langle\boldsymbol{\nu}_{j}\boldsymbol{\beta}^{T}\rangle\right]{\rm Tr}\left[{\bf u}_{j}{\bf b}^{T}\right]\right)+\ldots,$ (42)
The correlation terms on the last line of eq. (3.2) are model dependent, so additional assumptions are needed to proceed. We notice that in many cases the gradient vectors of ${\bf f}$ with respect to its $p+1$ parameters can be computed independently from each other. For example, this applies when $\mu$ and $\boldsymbol{\theta}$ are strength modifiers relative to templates obtained from independently simulated processes, which is, for example, the case considered by Barlow and Beeston in their original work BARLOW1993219 . Conversely, there is at least another relevant case in which the assumption of uncorrelated templates might seem inappropriate, that is when nuisance parameters are associated with systematic uncertainties estimated from event-by-event re-weightings of the “nominal” MC sample. In this case, correlations between the columns of ${\bf A}$ cannot be ignored a priori because of correlated Poisson fluctuations of the MC sample. However, if the dominant source of statistical uncertainty on the MC prediction originates from the variance of the event weights rather than from the purely Poisson uncertainty, different re-weightings of the same MC sample could be still treated as approximately independent. To proceed further, we make the assumption that the covariance matrix $\langle\boldsymbol{\nu}_{j}\boldsymbol{\beta}^{T}\rangle$ is zero and consider the remaining terms.
3.3 Contributions from the Jacobian matrix
Let us start by analyzing the ${\bf U}$-dependent bias terms appearing in eq. (3.2). The expectation values $\langle\boldsymbol{\nu}_{j}\rangle$ are generally different from zero. Indeed, from the condition $\lVert\hat{\bf u}_{j}\rVert=1$ it follows that
$02{\bf u}_{j}^{T}\langle\boldsymbol{\nu}_{j}\rangle=-\langle\boldsymbol{\nu}_{j}^{2}\rangle<$ 0 . (43)
@@ 3.3 Contributions from the Jacobian matrix
where ${\langle\boldsymbol{\alpha}_{j}^{2}\rangle}$ is the variance of the angle by which the vector ${\bf a}_{j}$ is rotated. It is easy to see that ${\langle\boldsymbol{\alpha}_{j}^{2}\rangle}$ is of order $n\cdot\Delta_{j}^{2}$, with $\Delta_{j}$ being the average relative MC uncertainty on the elements of ${\bf a}_{j}$. Under the same assumptions, one can easily prove that $\langle\boldsymbol{\alpha}_{j}\boldsymbol{\alpha}_{j}^{T}\rangle$ must be a matrix that projects onto the subspace orthogonal to ${\bf a}_{j}$ with equal probability along each of its $n-1$ directions, that is
@@ 3.3.1 Orthogonal column vectors
We start by considering the case ${\bf a}_{j}^{T}{\bf a}_{k}=\delta_{jk}$, that is the column vectors of the Jacobian matrix ${\bf A}$ are orthogonal to each other, which also implies a diagonal matrix ${\bf W}$. This hypothesis makes the analytic treatment simpler and its interpretation more intuitive, without precluding the general validity, as discussed later. It also has some physical interest: for example, it is appropriate when nuisance parameters parametrize systematic effects that behave in an almost “orthogonal” way in the space of observables ${\bf y}$, for example when they are determined from a PCA analysis Pearson01111901 ; Hotelling . Under the assumption of orthonormal column vectors, we can exploit the arbitrariness in the definition of the ${\bf u}_{j}$ basis by choosing ${\bf u}_{j}={\bf a}_{j}$. The second line on eq. (3.2) can be expanded by means of eq. (44) and (45), obtaining after some straightforward calculation:
$\frac{n}{n-1}\sum_{j=1}^{p}\left({\bf b}^{T}{\bf a}_{j}\right)^{2}\langle\boldsymbol{\alpha}_{i}^{2}\rangle-\frac{{\bf b}^{T}{\bf b}}{n-1}\left(\sum_{j=1}^{p}\langle\boldsymbol{\alpha}_{i}^{2}\rangle\right)$ (46)
@@
Let us first consider the case of large global correlation, that is we let $\rho_{\mu}\to 1$. Neglecting for a moment the numerical pre-factors, the two pieces appearing in eq. (46) are of the same order of magnitude, but of opposite sign. However, the former is scaled by a factor of $n$, thus giving a dominant positive contribution when the number of bins is large, which is our case of interest. This can be intuitively understood by noticing that when $n\gg p$, statistical fluctuations on ${\bf A}$ have a larger probability of shuffling around the basis vectors by introducing new directions along which ${\bf b}$ can have a finite projection, rather than rotate the vectors inside the same space, which would leave $S$ unchanged. These extra contributions add quadratically to $\lVert{\bf b}_{\perp}\rVert^{2}$ and can be numerically relevant if the latter is truly small. The same conclusion can be proved for generic $n$ under additional assumptions. Indeed, by expanding ${\bf b}$ as a linear combination of the vectors ${\bf a}_{j}$, and neglecting the ${\bf b}_{\perp}$ term, we have
@@
We now consider the other limiting case, that is $\rho_{\mu}\to 0$. The first term on the second line of eq. (46) can be now neglected compared to the second term, which is instead always negative. Thus, we see that in this case the effect of random fluctuations of the MC templates is to reduce the value of $\langle\hat{S}\rangle$ relative to ${S}$ by a factor of
$1-\frac{1}{n-1}\left(\sum_{j=1}^{p}\langle\boldsymbol{\alpha}_{i}^{2}\rangle\right)$ (48)
@@
This result is also in line with our intuition: if ${\bf b}$ is already orthogonal to the linear space spanned by the columns of ${\bf A}$, fluctuations on the latter can only generate spurious components of ${\bf b}$ in ${\rm Span}\{{\bf A}\}$, thus reducing the norm of ${\bf b}_{\perp}$. However, we notice that the relative decrease implied by eq. (48) is only of order of the squared relative uncertainty on the elements of $\tilde{\bf A}$, and is not scaled by $n$, so its overall impact might well be numerically less relevant than the effect discussed in the previous case.
Equations (46) and (48) can be verified, for example, in a simple two-dimensional setup, namely for $n=2$ and $p=1$. Modulo an overall normalization factor to the gradient vector and to the single-column Jacobian matrix, we can define
where $s_{\gamma}\equiv\sin\gamma$, $c_{\gamma}\equiv\cos\gamma$, with $\gamma\in[0,\pi]$ an angle that parametrizes the degree of correlation between $\mu$ and the nuisance parameter $\theta$, as illustrated by the left panel of figure 4. From eq. (49) we have that $S=U_{11}=c^{2}_{\gamma}$ and $\rho^{2}_{\mu}=s^{2}_{\gamma}=1-S$. We treat statistical fluctuations on ${\bf A}$ as a pair of independent and Gaussian-distributed random variables $(\alpha_{1},\alpha_{2})$, which, for simplicity, are assumed to have the same variance $\langle\alpha^{2}\rangle$. Under fluctuations, the matrix ${\bf U}$ is then changed to
Using eq. (50) evaluated at randomly generated points $(\alpha_{1},\alpha_{2})$, the estimator $\hat{S}=\hat{U}_{11}$ can be calculated and its mean value $\langle\hat{S}\rangle$ extracted from a large number of random samplings. The results are shown on the right of figure 4 as a function of the global correlation $\rho^{2}_{\mu}$ and for three representative values of $\langle\alpha^{2}\rangle$. For small values of $\rho^{2}_{\mu}$, the ratio $\langle\hat{S}\rangle/S$ is slightly less than one, i.e. the bias on $\hat{S}$ is negative, in agreement with eq. (48). Conversely, when $\rho^{2}_{\mu}$ increases, the ratio becomes larger than one (hence, the bias gets positive), eventually growing as $\langle\alpha^{2}\rangle/(1-\rho^{2}_{\mu})$ for $\rho^{2}_{\mu}\to 1$, in agreement with eq. (46).
Figure 4: Left: sketch of the two-dimenensional problem of eq. (49). Right: the ratio between the expectation value $\langle\hat{S}\rangle$ and the true value $S$ for the two-dimensional problem of eq. (49) as a function of the squared correlation coefficient $\rho^{2}_{\mu}$, and for three representative values of $\langle\alpha^{2}\rangle$.
3.3.2 Generic column vectors
Finally, we consider the more general case for the columns vectors ${\bf a}_{j}$. In this case, we can always find a set of $p$ orthonormal vectors such that
${\bf u}_{j}=\sum_{k=1}^{p}w_{jk}{\bf a}_{k}$ (51)
Again, we first focus on the case of large overlap between ${\bf b}$ and ${\rm Span}\{\bf A\}$, which allows us to consider only the first term on the right-hand side of eq. (52) since the remaining terms are relatively suppressed by a factor of $1/n$. If we denote by ${\bf v}$ a vector with $k^{\rm th}$-element equal to ${\bf b}^{T}{\bf a}_{k}$, we see that the triple sum of eq. (52) can be written as
@@ 3.3.2 Generic column vectors
For intermediate values of the global correlation, the existence of competition between positive and negative bias terms to $\langle\hat{S}\rangle$ does not allow for a conclusive statement on the overall sign, and the bias term should be determined case by case.
3.4 Contribution from the gradient
@@
Finally, we consider the role of statistical fluctuations on the gradient vector. Since the MC prediction for $\hat{\bf b}$ is randomly distributed around the true value ${\bf b}$, the expectation value of $\boldsymbol{\beta}$ is zero, so that terms linear in $\langle\boldsymbol{\beta}\rangle$ do not contribute to the bias term in eq. (3.2). The quadratic term on the first line of eq. (46) is positive-definite. This can be also seen as consequence of Jensen’s inequality applied to $\hat{S}$, which is a convex function of $\hat{\bf b}$. If we make the reasonable assumption $\langle\beta_{i}\beta_{j}\rangle=\langle\beta^{2}_{i}\rangle\delta_{ij}$, that is the MC predictions for different bins are statistically uncorrelated, the quadratic term can be further expanded as
@@
where the two terms in parentheses correspond to different ways of averaging the bin-by-bin variances $\langle\beta_{i}^{2}\rangle$. If the distribution of the latter is narrowly concentrated around an average value $\overline{\langle\beta^{2}\rangle}$, the right-hand side of eq. (54) is of order $\sim(n-p)\overline{\langle\beta^{2}\rangle}$, so it grows linearly with the number of bins. Thus, we see that statistical fluctuations on ${\bf b}$ can induce sizable positive biases to $\langle\hat{S}\rangle$ regardless of the global correlation coefficient, provided that $(n-p)\overline{\langle\beta^{2}\rangle}$ is comparable to $S$.
@@ 3.5 Discussion and connection with the toy model
We now summarize the assumptions and the main implications of eq. (3.2). We have assumed that the MC estimators for ${\bf b}$ and ${\bf A}$ are independent from each other and that statistical fluctuations of their elements can be treated as Gaussian-distributed noise around their true values. We have considered separately the impact of MC fluctuations on the Jacobian matrix ${\bf A}$ and on the gradient vector ${\bf b}$ in the form of a bias term to the quadratic form $S$, defined by eq. (32). If ${\bf b}$ has a large overlap with the linear space spanned by the columns of ${\bf A}$, or, equivalently, when the global correlation coefficient (38) is sizable, then the bias terms associated with fluctuations of the Jacobian matrix are positive-definite and increase with the number of bins $n$. Fluctuations of the gradient vector ${\bf b}$ also grow with $n$ and are always positive-definite. In both cases, $\hat{S}$ would differ from the true value $S$ by a positive offset, implying that the Hessian uncertainty, estimated as $\hat{\sigma}_{\rm H}=\hat{S}^{-1/2}$, would be underestimated [^9] More correctly, a positive bias on $\hat{S}$ would also imply a negative bias on $\hat{\sigma}_{\rm H}$ if the variance of $\hat{S}$ is small, so that $\langle\hat{S}^{-1/2}\rangle\approx\langle\hat{S}\rangle^{-1/2}$. In particular, by Taylor expansion, we find that $\frac{3}{8}\frac{\langle\hat{S}^{2}\rangle-\langle\hat{S}\rangle^{2}}{\langle\hat{S}\rangle^{2}}\ll 1$ should hold, a condition that is more likely to be satisfied when the number of bins is large, as for the law of large numbers..
@@
These findings can finally explain the observations made in the context of the toy model of sec. 2: when the original parametrization is considered (i.e. $\mu=\mu_{1}$), the linear correlation $\rho=\rho_{\mu}$ is indeed close to unity, implying a large bias term from fluctuations of the Jacobian matrix; for $\mu=\mu_{1}^{\prime}$, instead, the smaller correlation reduces the impact of a noisy Jacobian matrix, and the bulk of finite-size MC effects are now propagated to $\hat{\sigma}_{\rm H}$ through a noisy gradient vector $T_{1i}^{\prime}$.
@@
Remarkably, none of the bias terms appearing in eq. (46) is directly proportional to the true value $S$ by a naive factor of ${\cal O}(1)\cdot k^{-1}$, with $k$ related to the effective statistical power of the simulation relative to the data. The lack of proportionality also means that the bias and true value can stand in a ratio which is a priori arbitrary. This is one of the main observations of this work. Incidentally, this also explains how the Barlow-Beeston lite approach BBlite can be insufficient to account for MC uncertainties. As discussed in sec. 13, the lite version of the Barlow-Beeston likelihood is practically equivalent to a rescaling of the Poisson variance of the data by a factor of $(1+k^{-1})$, implying an increase of the Hessian uncertainty $\hat{\sigma}_{\rm H}$ by the square root of the same factor. However, a bias in $\hat{S}$, which is both positive and comparable to $S$, might reduce $\hat{\sigma}_{\rm H}$ by a potentially larger factor, ultimately leading to the non-physical situation of confidence intervals that are smaller than they would be in the limit of infinite-size MC samples.

### v3

@@ 3.2 Impact of MC fluctuations on $\sigma_{\rm H}$
The results presented in section 3.1.1 assume that the model $\bf f$} is known exactly. Suppose instead that $\bf f$} is determined from MC samples consisting of finite numbers of events, thus resulting in perturbed estimators $\hat{\bf b}={\bf b}+\boldsymbol{\beta}$ and $\hat{\bf A}$, with the latter giving rise to a perturbed matrix $\hat{\bf U}$, defined as in eq. (34) with the replacement $^{\bf A}\to\hat{\bf A}$. Since $\hat{\bf U}$ must also be idempotent, it admits an expansion in terms of a new basis of eigenvectors $\hat{\bf u}_{j}={\bf u}_{j}+\boldsymbol{\nu}_{j}$. By expanding the perturbed value $\hat{S}$ around its true value $S$, and taking the expectation value in the space of MC predictions, that is the mean value of $\hat{S}$ if infinitely many identical replicas of the MC samples could be generated, we obtain:
@@
$-2\sum_{j=1}^{p}\left({\rm Tr}\left[\langle\boldsymbol{\nu}_{j}\boldsymbol{\beta}^{T}\rangle\,{\bf u}_{j}{\bf b}^{T}\right]+{\rm Tr}\left[\langle\boldsymbol{\nu}_{j}\boldsymbol{\beta}^{T}\rangle\right]{\rm Tr}\left[{\bf u}_{j}{\bf b}^{T}\right]\right)+\ldots.$ (39)
The derivation of eq. (3.2) from a perturbative expansion of $\hat{S}$ is presented in D.2. Figure 4 helps visualize the geometrical interpretation of $^{\hat{S}}$, and the effect of statistical fluctuations on $\bf A$} and $\bf b$} for a representative problem with $n=2$.
Figure 4: A sketch showing the effect of statistical fluctuations on the relevant quantities introduced in the text. The simple case $n=2$ and $p=1$ has been considered.
The correlation terms on the last line of eq. (3.2) are model dependent, so additional assumptions are needed to proceed. We notice that in many cases the columns of the Jacobian of $\bf f$} with respect to its $p+1$ parameters can be computed independently from each other. For example, this applies when $\mu$ and $\boldsymbol{\theta}$ are strength modifiers relative to templates obtained from independently simulated processes, which is, for example, the case considered by Barlow and Beeston in their original work [3]. Conversely, there is at least another relevant case in which the assumption of uncorrelated templates might seem inappropriate, that is when nuisance parameters are associated with systematic uncertainties estimated from event-by-event re-weightings of the “nominal” MC sample. In this case, correlations between the columns of $\bf A$} cannot be ignored a priori because of correlated Poisson fluctuations of the MC sample. However, if the dominant source of statistical uncertainty on the MC prediction originates from the variance of the event weights rather than from the purely Poisson uncertainty, different re-weightings of the same MC sample could be still treated as approximately independent. To proceed further, we make the assumption that the covariance matrix $\langle\boldsymbol{\nu}_{j}\boldsymbol{\beta}^{T}\rangle$ is zero and consider the remaining terms in eq. (3.2).
3.3 Contribution from the $\bf A$} matrix
Let us start by analyzing the bias terms in eq. (3.2) related to fluctuations of the $\bf A$} matrix. The expectation values $\langle\boldsymbol{\nu}_{j}\rangle$ are generally different from zero. Indeed, from the normalization condition $\lVert\hat{\bf u}_{j}\rVert=1$ it follows that
$2{\bf u}_{j}^{T}\langle\boldsymbol{\nu}_{j}\rangle=-\langle\boldsymbol{\nu}_{j}^{2}\rangle\neq 0.$ (40)
@@ 3.3 Contribution from the $\bf A$} matrix
where $\langle\boldsymbol{\alpha}_{j}^{2}\rangle$} is the variance of the angle by which the vector ${\bf a}_{j}$ is rotated. It is easy to see that $\langle\boldsymbol{\alpha}_{j}^{2}\rangle$} must be of order $n\cdot\Delta_{j}^{2}$, with $\Delta_{j}$ being the average relative MC uncertainty on the elements of ${\bf a}_{j}$. Under the same assumptions, one can easily prove that $\langle\boldsymbol{\alpha}_{j}\boldsymbol{\alpha}_{j}^{T}\rangle$ must be a matrix that projects onto the subspace orthogonal to ${\bf a}_{j}$ with equal probability along each of its n1n-1 directions, that is
@@
In the following, we will restrict to the case ${\bf a}_{j}^{T}{\bf a}_{k}=\delta_{jk}$, that is the column vectors of the matrix $\bf A$} are orthogonal to each other. This hypothesis makes the analytic treatment simpler and its interpretation more intuitive, without precluding the general validity, as discussed in D.3. We also notice that this special case has some physical interest: for example, it is appropriate when nuisance parameters parametrize systematic effects that behave in an almost “orthogonal” way in the space of observables $\bf y$}, for example when they are determined from a PCA analysis [22, 18].
Under the assumption of orthonormal column vectors, we can exploit the arbitrariness in the definition of the ${\bf u}_{j}$ vectors by choosing ${\bf u}_{j}={\bf a}_{j}$. The second line on eq. (3.2) can be expanded by means of eq. (41) and (42), obtaining after some straightforward calculation:
$\frac{n}{n-1}\sum_{j=1}^{p}\left({\bf b}^{T}{\bf a}_{j}\right)^{2}\langle\boldsymbol{\alpha}_{i}^{2}\rangle-\frac{{\bf b}^{T}{\bf b}}{n-1}\left(\sum_{j=1}^{p}\langle\boldsymbol{\alpha}_{i}^{2}\rangle\right).$ (43)
@@
Let us first consider the case of a large value of $\rho_{\mu}$. Aside from numerical factors, the two pieces appearing in eq. (43) are of the same order of magnitude, but of opposite sign. However, the former is scaled by a factor of $n$, thus giving a dominant positive contribution when the number of bins is large, which is our case of interest. This can be intuitively understood by noticing that when $n\gg p$, statistical fluctuations on $\bf A$} have a larger probability of shuffling around the basis vectors by introducing new directions along which $\bf b$} can have a finite projection, rather than rotate the vectors inside the same space, which would leave $S$ unchanged. These extra contributions add quadratically to $\lVert{\bf b}_{\perp}\rVert^{2}$ and can be numerically relevant if the latter is small. As discussed in D.4, the same conclusion can be proved for generic $n$ under some more restrictive assumptions.
@@
We now consider the opposite limiting case, that is $\rho_{\mu}\to 0$. The first term on the second line of eq. (43) can be now neglected compared to the second term, which is instead always negative. Thus, we see that in this case the effect of random fluctuations of the MC templates is to reduce the value of $\langle\hat{S}\rangle$ relative to S{S} by a factor of
$1-\frac{1}{n-1}\left(\sum_{j=1}^{p}\langle\boldsymbol{\alpha}_{i}^{2}\rangle\right).$ (44)
@@
This result is also in line with our intuition: if $\bf b$} is already orthogonal to the linear space spanned by the columns of $\bf A$}, fluctuations on the latter can only generate spurious components of $\bf b$} inside the linear space ${\cal V}_{\bf A}$, thus reducing the norm of $\bf b$$_{\perp}$. However, we notice that the relative decrease implied by eq. (44) is only of order of the squared relative uncertainty on the elements of $\tilde{\bf A}$, and is not scaled by $n$, so its overall impact might well be numerically less relevant than the effect discussed in the previous case.
@@
For intermediate values of $\rho_{\mu}$, competition between positive and negative bias terms to $\langle\hat{S}\rangle$ does not allow for a conclusive statement on the overall sign, and the bias term should be determined case by case. This can be well illustrated, for example, in a two-dimensional setting, as discussed in detail in D.5. In this simplified setup, the ratio $\langle{\hat{S}}\rangle/S$ is found to depend on $\rho_{\mu}$ as expected from eqs. (43) and (44).
3.4 Contribution from the $\bf b$} vector
@@
Finally, we consider the role of statistical fluctuations on the $\bf b$} vector. Since the MC prediction for $\hat{\bf b}$ is randomly distributed around the true value $\bf b$}, the expectation value of $\boldsymbol{\beta}$ is zero, so that terms linear in $\langle\boldsymbol{\beta}\rangle$ do not contribute to the bias term in eq. (3.2). The quadratic term on the first line of eq. (43) is positive-definite [^9] This can be also seen as consequence of Jensen’s inequality applied to $\hat{S}$, which is a convex function of $\bf b$}.. If we make the reasonable assumption $\langle\beta_{i}\beta_{j}\rangle=\langle\beta^{2}_{i}\rangle\delta_{ij}$, that is the MC predictions for different bins are statistically uncorrelated, the quadratic term can be further expanded as
@@
where the two terms in parentheses correspond to different ways of averaging the bin-by-bin variances $\langle\beta_{i}^{2}\rangle$. If the distribution of the latter is narrowly concentrated around an average value $\overline{\langle\beta^{2}\rangle}$, the right-hand side of eq. (45) is of order $\sim(n-p)\overline{\langle\beta^{2}\rangle}$, so it grows linearly with the number of bins. Thus, we see that statistical fluctuations on $\bf b$} can induce sizable positive biases to $\langle\hat{S}\rangle$ for any value of $\rho_{\mu}$, provided that $(n-p)\overline{\langle\beta^{2}\rangle}$ is comparable to $S$.
@@ 3.5 Discussion and connection with the toy model
We now summarize the assumptions and the main implications of eq. (3.2). We have assumed that the MC estimators for $\bf b$} and $\bf A$} are independent from each other and that statistical fluctuations of their elements can be treated as Gaussian-distributed noise around their true values. We have considered separately the impact of MC fluctuations on the blocks $\bf A$} and $\bf b$} of the Jacobian matrix in the form of a bias term to the quadratic form $S$, defined by eq. (35). If $\bf b$} has a large overlap with the linear space spanned by the columns of $\bf A$}, or, equivalently, when $\rho_{\mu}$ is close to one, then the bias terms associated with fluctuations of the $\bf A$} matrix are positive-definite and increase with the number of bins $n$. Fluctuations of the $\bf b$} vector also grow with $n$ and are always positive-definite. In both cases, $\hat{S}$ would differ from the true value $S$ by a positive offset, implying that the Hessian uncertainty, estimated as $\hat{\sigma}_{\rm H}=\hat{S}^{-1/2}$, would be underestimated [^10] More correctly, a positive bias on $\hat{S}$ would also imply a negative bias on $\hat{\sigma}_{\rm H}$ if the variance of $\hat{S}$ is small, so that $\langle\hat{S}^{-1/2}\rangle\approx\langle\hat{S}\rangle^{-1/2}$. In particular, by Taylor expansion, we find that $\frac{3}{8}\frac{\langle\hat{S}^{2}\rangle-\langle\hat{S}\rangle^{2}}{\langle\hat{S}\rangle^{2}}\ll 1$ should hold, a condition that is more likely to be satisfied when the number of bins is large, as for the law of large numbers..
@@
These findings can finally explain the observations made in the context of the toy model of sec. 2: when the original parametrization is considered (i.e. $\mu=\mu_{1}$), the linear correlation coefficient $\rho=\rho_{\mu}$ is indeed close to unity, implying a large bias term from fluctuations of the matrix ${\bf A}=(T_{21}/\sqrt{y_{1}},\ldots,T_{2n}/\sqrt{y_{n}})^{T}$; for $\mu=\mu_{1}^{\prime}$, instead, the smaller correlation reduces the impact of statistical fluctuations on $\bf A$}, and the bulk of finite-size MC effects are now propagated to $\hat{\sigma}_{\rm H}$ through a noisy vector ${\bf b}=(T_{11}^{\prime}/\sqrt{y_{1}},\ldots,T_{1n}^{\prime}/\sqrt{y_{n}})^{T}$.
@@
Remarkably, none of the bias terms appearing in eq. (43) is directly proportional to the true value $S$ by a naive factor of ${\cal O}(1)\cdot k^{-1}$, with $k$ related to the effective statistical power of the simulation relative to the data. The lack of proportionality also means that the bias and true value can stand in a ratio which is a priori arbitrary. This is one of the main observations of this work. Incidentally, this also explains how the Barlow-Beeston lite approach [7] can be insufficient to account for MC uncertainties. As discussed in sec. 16, the lite version of the Barlow-Beeston likelihood is practically equivalent to a rescaling of the Poisson variance of the data by a factor of $(1+k^{-1})$, implying an increase of the Hessian uncertainty $\hat{\sigma}_{\rm H}$ by the square root of the same factor. However, a bias in $\hat{S}$, which is both positive and comparable to $S$, might reduce $\hat{\sigma}_{\rm H}$ by a potentially larger factor, ultimately leading to the non-physical situation of confidence intervals that are smaller than they would be in the limit of infinite-size MC samples.